A company has sales of $2.5m, a gross margin percent of 40% and a net income of $500,000. The balance sheet shows total assets of $1.1m; total liabilities of $600,000; retained earnings of $350,000 and notes payable of $150,000.

Please calculate the following:

a.   Return on Equity:
b.   Profit margin:
c.   Return on Assets:
d.   Cost of goods sold:
Answer #1

a.

ROE = 500,000/350,000 = 142.86%

b.

Profit Margin = 500,000/2,500,000 =20.%

c.

ROA = 500,000/(600,000 + 350,000 + 150,000) = 45.45%

d.

COGS = 0.60(2,500,000) = $1,500,000